Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does X receive than if the profits were divided in proportion to the amount invested?
: Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does X receive than if the profits were divided in proportion to the amount invested?

Answer by stanbon(18999)

Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does X receive than if the profits were divided in proportion to the amount invested?
---------
Equal amount received by each 4800/3 = $1600 for each
--------------------------
Proportional amount received by each:
Proportion: 9:7:4
9x + 7x + 4x = 4800
20x = 4800
x = 240
x amount: 9x = 2160
y amount: 7x = 1680
z amount: 4x = 960
---------------------
how much less does X receive than if the profits were divided in proportion to the amount invested?
2160 - 1600 = $560
=======================
Cheers,
Stan H.